substitution method
2x-y=-5
x+5y=14

Answer by rmromero(383)


substitution method
2x-y=-5
x+5y=14

Here are the Steps using Substition method
1. In either equation, solve for one variable in terms of the other.
2x - y = -5
-y = -2x - 5
y = 2x + 5

2. Substitute for that variable in the other equation. Solve.
x + 5y = 14
x + 5(2x+5) = 14
x + 10x + 25 = 14
11x + 25 = 14
11x = -11
x = -1


3. Substitute the result from step 2 in either equation. Solve for the other variable.
2x - y = -5, x = -1
2(-1) - y = -5
-2 - y = -5
-y = -3
y = 3

4. Check the solution in both original equations.
x = -1 and y = 3
2x - y = -5
2(-1) - (3) = -5
-2 - 3 = -5
-5 = -5 ----> True


x + 5y = 14
(-1) + 5(3) = 14
-1 + 15 = 14
14 = 14 --------> true

Therefore the solution is (-1, 3)